OpenStudy (anonymous):

If x^3=y^9, what is x in terms of y? √(y) y^2 y^3 y^6 y^12

OpenStudy (mathmate):

Hint: y^9=(y^3)^3

OpenStudy (anonymous):

here i need to solve for x

OpenStudy (mathmate):

x^3=y^9=(y^3)^3 what can you say about x?

OpenStudy (anonymous):

\[(x)^3=(y^3)^3\]

OpenStudy (anonymous):

@henpen so its y^9?

OpenStudy (anonymous):

Not quite. x^3 is y^9, so what is x (hint: x is the cube root of x^3)?

OpenStudy (anonymous):

idk

OpenStudy (mathmate):

If \[a ^{3}\ = b^{3}\] what can you say about a and b?

OpenStudy (anonymous):

what can I say? u mean, what the values of a and b are?

OpenStudy (mathmate):

We don't know a nor b, but there is a relationship between a and b we can say. For example, x^3=4^3, what can you say about x?

OpenStudy (anonymous):

\[x^3=4^3\] Means: there is some number, that when you cube it it spits out the same number as if you were to cube 4.

OpenStudy (anonymous):

right

OpenStudy (mathmate):

So what is the value of x in this case?

OpenStudy (anonymous):

y^3?

OpenStudy (mathmate):

Is you answer for the original question or the example x^3=4^3?

OpenStudy (anonymous):

question

OpenStudy (mathmate):

Good job! Is everything clear how you would go about solving a similar problem, such as: y^5=x^15, Can you express y in terms of x?

OpenStudy (anonymous):

yes. thnx a lot!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

y^5=x^15 x=y^(1/3)

OpenStudy (mathmate):

That is correct for x in terms of y. What about y in terms of x?

OpenStudy (anonymous):

y= x^3

OpenStudy (mathmate):

Perfect! I can see that you understand the concept!

OpenStudy (anonymous):

I didnt understand it at first when i was asking the question :)

OpenStudy (mathmate):

That's why we're all here, to get things cleared up.